WLCG-OSG-EGEE Ops' Minutes Mon 19 Jan 2009


For LHCb, CNAF is currently unusable, so will be dropped from the FEST09 exercise.


Please remember to connect to the web conference even if you dial-in, so that we don't have to guess who's out there! e.g. sip:anonymous is not particularly helpful!


  • Asia Pacific ROC: absent
  • Central Europe ROC: Malgorzata Krakowian
  • OCC / CERN ROC: Maite Barroso, John Shade, Antonio Retico
  • French ROC: 0033478930880
  • German/Swiss ROC: Angela Poschlad
  • Italian ROC: Alessandro Paolini
  • Northern Europe ROC: Ron Trompert
  • Russian ROC: Lev Shamardin, Victor Edneral
  • South East Europe ROC: Kostas Koumantaros
  • South West Europe ROC: Kai Neuffer
  • UK/Ireland ROC: Derek Ross
  • representing C-COD: Vera Hansper
  • GGUS: Helmut Dres


  • WLCG Service Cordination: Harry Renshall

WLCG Tier 1 Sites

  • ASGC: absent
  • BNL: Absent
  • CERN site: Absent
  • FNAL: Absent
  • FZK: Angela Poschlad
  • IN2P3: 0033478930880
  • INFN: Alessandro Paolini
  • NDGF: Vera Hansper
  • PIC: Kai Neuffer
  • RAL: Gareth Smith, Derek Ross
  • SARA/NIKHEF: Absent
  • TRIUMF: Absent

LHC Experiments

  • ATLAS: absent (Alessandro conveyed his apologies)
  • LHCb: Roberto Santinelli
  • CMS: absent
  • ALICE: absent

Special guests

  • Di Qing, Remi Mollon, Laurence Field (CERN)

Feedback on Last Week's Minutes

None was given.

EGEE Items

Grid Operator Hand Over on Duty

  Primary Team Secondary Team
From ROC UKI ROC Italy

  • BDII issues raised by three ROCs via CODs. Will be discussed later on.
  • ENEA-INFO has sprung into life so no escalation is needed. Gareth will reset to 1st step.
  • Alessandro from Italy had nothing to add for grid operator handover.

Antonio hadn't arrived yet to give his gLite release news, so we jumped to issues from ROC reports, namely the BDII update 34 problems. Laurence had nothing to add. Yet again, the conference system played up, so we rebooted the phone & tried again. These recurring phone problems are a big waste of everyone's time!

EGEE Items From ROC Reports

  • BDII: new release is in certification (to be installed at your own risk). Antonio said plan for release was to have CNAF, GRIF & AEGIS to keep testing it. Kostas wanted a release date. Antonio: certification this week, wait another week in Production, and then release.
  • Another issue from Italy with thousands of defunct Globus-gma jobs. Di Qing has released a patch to clean up defunct processes (done this afternoon). Link is in GGUS:42981 for those who want it quickly (patch is being certified). Italy will try out the patch and give feedback.
  • APEL accounting data to GOC was broken. Spotted by SAM tests and solved. The "too long time to solve" was due to a lack of updates to GGUS ticket by different support entities.

Maite read from the agenda about the complaint concerning the stability of the BDII (ever-increasing functionality but no improved reliability). Laurence commented that the Savannah bug was not about BDII stability, but rather robustness of the Information System (i.e. GFAL failover to different BDIIs). Rémi stated that GFAL (as of GFAL already has BDII failover mechanism for when BDII is down, but that empty responses (i.e. comatose BDII) were harder to handle.

PPS Reports and Issues

As usual, please refer to the link in the agenda for the detailed information!
  • 3.1.100 WMS release is about to be certified. Pilot service will be set up, so Antonio requests a Prod site to deploy the service as soon as the patch is certified, and have experiments use it. Volunteers to contact Antonio.
  • CREAM pilot has yet another new version, available in repository, for sites participating.
  • New version of VOMS client with bug-fixes is being certified
  • Antonio anticipates a (phased) move to production of all the above sometime in February.

gLite Release News

As usual, please refer to the link in the agenda for the detailed information!
  • gLite 3.1 update 39 in PPS. New version 1.6.1 of VDT available. Used by everyone, but a rollback procedure needed before a roll-out can be considered.

WLCG Items

WLCG issues coming from ROC reports

  • None

Upcoming WLCG Service Interventions

  • Consult links on the agenda page.

WLCG Service Coordination

  • Harry mentioned that the "ATLAS 10 million file test" is now scheduled to run until Sunday (there were CERN FTS problems due to the non-renewal of a proxy on an FTS server).

ATLAS Service

ALICE Service

No report and no one present, but Harry mentioned that they’re setting up a WMS in GRIF this week.

CMS Service

CMS report is on the agenda page.

LHCb Service

Roberto delved into the long-standing issue that LHCb has with the CNAF shared area:
  • increasing timeout for setting up environment still doesn’t solve CNAF’s shared area problem. (CNAF will buy new per-VO hardware in 2 months, but not fast enough for LHCb). CNAF currently unusable, so will be dropped from FEST09 exercise.
  • CNAF GGUS:44729 ticket opened prior to Christmas, but closed/re-opened/closed/re-opened several times. Complaint about closing tickets without end-user approval!
  • native 64-bit O/S still had 32-bit library problems at PIC & CNAF. Extra libraries were needed, so this has been added explicitly to the LHCb VO card. The same problem of missing packages is present in SL5 frown

OSG Items

No one from OSG was on the line, but Maria, for the record, complained that there were three tickets untouched since November:

Newly Created Action Items

Assigned to Due date Description State Closed Notify  
Main.OCC 2007-03-05 Example Action Item 2007-03-06 SteveTraylen   edit

Review of Open Action Items

Open Action Items

IdSubmitterDescriptionCreationDueAssigned To 

Actions Closed in Last 20 Days

IdSubmitterDescriptionCreationDueAssigned ToClosed 


None was raised.

Next Meeting

The next meeting will be Monday, 26 Jan 2009 15:00 UTC (16:00 Swiss local time).

  • Attendees can join from 14:45 UTC (15:45 Swiss local time) onwards.
  • The meeting will start promptly at 15:00 UTC (16:00 Swiss local time).
  • The WLCG section will start at the fixed time of 15:30 UTC (16:30 Swiss local time).
  • To dial in to the conference:
    • Dial +41227676000
    • Enter access code 0148141

These minutes can only be changed by members of:

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2009-01-20 - JohnShade
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    EGEE All webs login

This site is powered by the TWiki collaboration platform Powered by Perl This site is powered by the TWiki collaboration platformCopyright &© by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Ask a support question or Send feedback